View Help Grammarly Tell me what you want to do to edit, its safer to stay in Protected View. Enable Editing 1. Fifty pro-football rookies were rated on a scale of 1 to 5 based on performance at a training camp. A ranking of 1 is a poor prospect and a ranking of 5 is an excellent prospect. Enter the below data into Excel. Construct a relative frequency and a cumulative relative frequency. Rating Frequency 10 14 18 a. How many of the rookies received a rating of 4 or better? b. How many of the rookies received a rating of 2 or worse? c. What percent received a rating of 5? d. What percent received a rating of 3 or better? A recent survey asked 5,000 people: Whats most important to you when choosing where to live? The responses are shown below. Enter the data into Excel. 2. Relative ResponseFrequency Good Jobs Affordable Homes 0.25 0.15 0.35 0.15 Schools Low Crime Things to do a. Construct the frequency distribution. You have to go backwards using 5,000 people. b. How many of the people chose good jobs as the most important? 1 The file (Math SAT) can be found in Blackboard. The file contains data on average math SAT scores for each state in the US. a. Construct a frequency distribution (relative & cumulative) and histogram using the data 3. analysis toolbar. You should use the 4 steps of constructing a histogram. Make the widths a nice round number Interpret the relative frequency b.
Answer #1
Rating Frequency
1 4
2 10
3 14
4 18
5 4
Total 50

a)

4 or better = 18+4 =22

b)

2 or worse = 10+4 =14

c) Percentage of rating 5 =

4 * 100 8.00%

d) Percentage of rating 3 or better =

1418 +4) 100 -O 72% 50